Suppose that Goldstar Bank is completely "loaned up." Now suppose that a customer deposits an additional $45,000 into the bank. Assume the reserve requirement is 20 percent. Instructions: Enter your answers as whole numbers. a. As a consequence of

Suppose that Goldstar Bank is completely "loaned up." Now suppose that a customer deposits an additional $45,000 into the bank. Assume the reserve requirement is 20 percent.

 

Instructions: Enter your answers as whole numbers.

 

a. As a consequence of the $45,000 deposit, Goldstar Bank will now have excess reserves in the amount of $  .

 

b. As a result of the $45,000 initial deposit into Goldstar Bank, the banking system can generate a maximum of $  in new deposits and loans.
Answer #1

Excess reserve is the reserve which a bank keeps after keeping the required reserve aside from which it lends.

If $45,000 has been deposited then 20% of its hould be kept aside as reserved from which loan can not be granted.

Thus excessreserve will be 45000 - 20% of 45000

Excess reserve = 45000 - 9000 = $36,000

Now all of this excess resreve can be given into loan. Thus now the bank can give 36000 to loan which will actually gets deposited into another bank and the another bank will now keep 20% of 36000 and at maximum can give rest into loans. hence the process will go on. Thus the banking system can generate maximum of 36000 - 20% of 36000 = $28800 into new deposits and loans.